I need step by step instructions on how to solve the following:
Solve by the elimination method.
9x - y = 40
x + 8y = 45

Answer by mananth(16946)

9 x + -1 y = 40 .............1
1 x + 8 y = 45 .............2
Eliminate y
multiply (1)by 8
Multiply (2) by 1
72 x -8 y = 320
1 x + 8 y = 45
Add the two equations
73 x = 365
/ 73
x = 5
plug value of x in (1)
9 x + -1 y = 40
45 + -1 y = 40
-1 y = -5

y = 5
(5,5)